# HidePrint: Hiding the Radio Fingerprint via Random Noise

### v1

@@ Abstract
Radio Frequency Fingerprinting (RFF) techniques allow a receiver to authenticate a transmitter by analyzing the physical layer of the radio spectrum. Although the vast majority of scientific contributions focus on improving the performance of RFF considering different parameters and scenarios, in this work, we consider RFF as an attack vector to identify and track a target device.
@@
We propose, implement, and evaluate HidePrint, a solution to prevent tracking through RFF without affecting the quality of the communication link between the transmitter and the receiver. HidePrint hides the transmitter’s fingerprint against an illegitimate eavesdropper by injecting controlled noise in the transmitted signal. We evaluate our solution against state-of-the-art image-based RFF techniques considering different adversarial models, different communication links (wired and wireless), and different configurations. Our results show that the injection of a Gaussian noise pattern with a standard deviation of (at least) 0.02 prevents device fingerprinting in all the considered scenarios, thus making the performance of the identification process indistinguishable from the random guess while affecting the Signal-to-Noise Ratio (SNR) of the received signal by only 0.1 dB. Moreover, we introduce the selective radio fingerprint disclosure, a new technique that allows the transmitter to disclose the radio fingerprint to only a subset of intended receivers. This technique allows the transmitter to regain anonymity, thus preventing identification and tracking while allowing authorized receivers to authenticate the transmitter without affecting the quality of the transmitted signal.
1 Introduction
@@
Radio Frequency Fingerprinting (RFF) is a popular technique used to differentiate and identify Radio Frequency (RF)-enabled devices by taking advantage of the unique characteristics inherent in the RF emissions [1]. These unique characteristics come from the hardware of the device, and in particular, by the small differences featured by the hardware components during the manufacturing process. Such differences lead to the main idea that two electronic devices that are exactly alike do not exist because of differences in oscillators, amplifiers, and built-in antennas. Since these differences are reflected in the over-the-air signal, a receiver can extract them through the analysis of the received signals and leverage them to uniquely identify the transmitter [2]. A strict requirement for all RFF systems is the ability to collect information from the physical layer of the communication link [3]. Such features persist on the receiver side up to the conversion between symbols to bits, making the availability of such information a requirement for RFF. At the physical layer, the baseband signal can be represented by In-Phase - Quadrature (IQ) samples representing the in-phase (I) and quadrature (Q) components, used to carry the information associated with the transmitted bit string and some additional information such as noise, attenuation, Doppler shift, and finally, the radio fingerprint [4].
@@
RFF features several key advantages compared to traditional cryptographic solutions. RFF is crypto less, i.e., it requires no implementation of cryptographic techniques, and, most importantly, it requires no key distribution, thus solving by design the challenges related to joining and leaving entities from the network. Moreover, RFF does not require any intervention on the transmitter side. The transmitter is authenticated by the information (not so) hidden) in its transmitted signal. In contrast, RFF requires a receiver capable of collecting information from the physical layer and a Deep Learning (DL) model. The receiver should train a neural network with legitimate signals (IQ samples) in order to challenge it in the subsequent phase, with a test set (of IQ samples) to assess either the presence or the absence of the transmitter(s) [5].
@@
Scientific literature in the latest years mostly focused on maximizing the performance of RFF techniques while addressing several real-world challenges, such as robustness to multipath fading [6], time-varying distortions [7], reliability over time [8], and mobility of the transmitter [9], to name a few. As a result, RFF solutions today are becoming increasingly effective for successfully classifying and authenticating the transmitter, making RFF a real candidate to implement authentication at the physical layer while mitigating the common challenges of key distribution in dense networks and implementation in resource-constrained devices. While a ready-to-market RFF solution is not yet available, the accuracy of the proposed techniques is increasing for the most diverse scenarios and applications [2].
@@
In this work, we propose a different perspective on RFF, exploring the possibility of adopting RFF by a malicious third party to detect and track the presence of a device in the radio spectrum. Indeed, an adversary could train a neural network model on signals generated by a specific (pool of) device(s) and then leverage such a model to detect the presence of the transmitter in the radio spectrum—making useless any anonymization technique implemented by the higher communications layers, e.g., MAC address anonymization. Malicious detection and tracking of devices pose significant challenges to people’s privacy, enabling mass and targeted surveillance. To provide an example, a practical use case might involve tracking when and where a device appears in the radio spectrum to infer a person’s daily routine, places of interest, and social interactions. Research on RFF-based attacks is relatively scarce in the literature and is mostly qualitative. Although pioneering contributions consider RFF techniques based on Channel State Information (CSI) fingerprinting [10, 11], a few contributions explored the usage of Adversarial Machine Learning (AML) techniques to generate perturbations in the transmitted signals thwarting RFF while keeping the communication quality acceptable [12][13][14], and [15]. However, such works either focus on pure data analysis without considering the technical challenges of data manipulation and wireless transmission or use simplistic RFF models, being far-away from real-world deployments. As a result, no contributions investigated the potential of intentionally altering the radio fingerprint to prevent the tracking of the transmitter while allowing authorized receivers to authenticate it without affecting the quality of the communication link.
@@
RQ1: Is it possible to prevent a malicious adversary from successfully using RFF to identify and track a (pool of) device(s)?
RQ2: Is it possible to design a selective radio fingerprint disclosure technique?
Since RFF does not require shared secrets, the legitimate and malicious receivers leverage the same information. Thus, a solution preventing RFF-based attacks by an adversary (RQ1) will affect the ability of a legitimate device to authenticate the transmitter(s), thus yielding to our second research question (RQ2). Selective radio fingerprint disclosure could allow a designated set of (legitimate) receivers to perform RFF while excluding the other (malicious) ones.
Contribution. This work provides manifold contributions to the RFF research domain, as listed below.
  • We design, implement, evaluate, and test two RFF-based attacks in different scenarios, considering both an (ideal) wired and a (real) wireless communication link.
@@
  • We consider an adversary capable of collecting data for RFF both before and after anonymization, so updating at runtime the RFF models to de-anonymize the transmitters.
@@
  • Inspired by HidePrint, we propose a technique enabling selective radio fingerprint disclosure, thus allowing (only) a selected subset of (legitimate) receivers to perform RFF while excluding the malicious ones.
Paper organization. The paper is organized as follows. Sect. 2 reviews related work about RFF, Sect. 3 introduces the adversary models and scenario assumptions used in our paper, Sect. 4 introduces the theoretical framework and rationale behind our solution, Sect. 5 provides the details of our measurement campaign and the considered hardware and software setup, Sect. 6 introduces our data pre-processing methodologies and DL techniques, Sect. 7 discusses the performance of our solution in both the wired and wireless scenarios, Sect. 8 evaluates our solution while considering competing RFF techniques, Sect. 9 presents the concept of selective radio fingerprint disclosure, Sect 10 discusses the main results and limitations of our solution, and finally, Sect. 11 concludes the paper and outlines future work.
2 Related Work
RFF received a lot of attention in recent years as a technique to effectively authenticate the transmitter device [1], [2].
@@ 2 Related Work
Radio Fingerprint
Disclosure
[10] CSI-based Phase Correction Real-world tests
[11] CSI-based Random Phase Correction Real-world tests
[12] Image-based AML Simulations
[13] Raw IQ AML Simulations
[14] Raw IQ AML Simulations
[15] Raw IQ AML Real-world tests
[16] Image-based Friendly Jamming Real-world tests
@@
Table 1: Qualitative comparison of scientific contributions focusing on RFF disruption.
@@
Early research on RFF investigated the usage of specific radio signal imperfections to identify transmitters and reject radio spoofing attacks, e.g., IQ imbalance, phase shift, and Carrier Frequency Offset (CFO), mostly using Machine Learning (ML) models [17][18][19]. With the diffusion of DL, research on RFF shifted towards using raw physical-layer signals, i.e., IQ samples, due to the enhanced capability of such algorithms to recognize patterns hidden in the noise [1]. Most of the solutions proposed until a few years ago leveraged raw IQ samples directly as input to various DL classifiers, e.g., Convolutional Neural Networkss [20], Long Short-Term Memorys [21], and Autoencoderss [9], to name a few. On the one hand, such solutions are characterized by outstanding classification and authentication performance, reaching accuracy values very close to perfect discrimination [3]. On the other hand, recent research showed that the models created through these approaches are very unstable, being significantly affected by any source of noise, introduced, e.g., by the warm-up of the radio [22], hardware reset of the radio [8], temperature changes in the environment [23] [24], fluctuations in channel conditions, and firmware reload [25]. In this context, image-based RFF techniques have emerged as a branch of RFF characterized by enhanced flexibility and portability across different real-world scenarios and configurations. Recent literature has shown that pre-processing (and converting) raw IQ samples into images leads to the creation of DL-based models robust to AML [12], restart of the radios [8], and firmware reload [25], to name a few. Image-based RFF techniques have also been used beyond device identification, e.g., for distinguishing between communication channels, i.e., channel fingerprinting [26, 27], and jamming detection [28, 29]
@@
Only a few works in the literature considered RFF as an attack and, thus, investigated how to protect a device from unconditionally disclosing the radio fingerprint to malicious parties. An early attempt in this direction is the contribution by Abanto et al. [11], showing that the injection of random phase corrections into the signal emitted by the devices can reduce the accuracy of RFF systems using CSI to authenticate a device. The RFF algorithm proposed in their paper only considers CFO and phase shift estimation, while DL-based solutions are not considered. Also, the pioneering contribution by Danev et al. in [10] considers CSI-based RFF and experiments with impersonation attacks via powerful signal generators. More recently, contributions such as the ones by Papangelo et al. [12], Sun et al. [14], and Liu et al. [13] investigated the effectiveness of AML to disrupt the correct operation of RFF. The only recent work considering the usage of adversarial techniques to conceal radio fingerprints is the preliminary work by Lu et al. in [15], providing an early solution to anonymization against RFF. The authors propose to generate a perturbation signal $\delta$ calculated according to a predefined strategy that matches the decision algorithm taken by the classifier. The adopted CNN network, designed from scratch, uses raw IQ samples, sharing the same limitations described above. Moreover, the authors use a limited set of devices (five) of different brands, easing the design of the RFF model and the following anonymization task, and use higher-order modulation schemes, being increasingly sensitive to even low-power noise. An alternative approach is the one proposed by Irfan et al. [16], using an external device (jammer) synchronized with the legitimate pool of devices to protect against RFF. Although effective, such an approach requires an external device, making the design of the solution more complex and enlarging the resulting threat surface.
Thus, as summarized in Tab. 1, no work in the literature investigated the ability to protect from malicious RFF while considering various adversary models, characterized by increasing knowledge of the RFF model of the target device(s), and state-of-the-art image-based RFF techniques, while avoiding to resort to additional devices. Moreover, there is currently no technique providing selective radio fingerprint disclosure, i.e., allowing the disclosure of the radio fingerprint only to selected receivers. We address this gap in the remainder of this paper.
3 Scenario and Adversarial model
@@
Our scenario is constituted by a pool of ten (10) transmitters that aim to remain anonymous while communicating. Current state-of-the-art techniques already allow to enforce anonymity at higher layers of the ISO OSI communication stack, e.g., at the network layer via VPN and at the MAC layer via MAC randomization, while our scenario specifically considers anonymity against physical layer identification. Without loss of generality, we consider the usage of the Binary Phase Shift Keying (BPSK) modulation scheme. Many modern communication technologies currently use BPSK, including IEEE 802.15.4 (Zigbee, Thread), IEEE 802.11 (WiFi), and many satellite communication systems (e.g., IRIDIUM) due to its relevant robustness to noise introduced by the wireless channel [30]. Moreover, we consider an adversary able to deploy state-of-the-art image-based RFF techniques to identify and (potentially) track devices in the radio spectrum. We recall that, as discussed in Sec. 2, image-based RFF models typically outperform other solutions thanks to their enhanced robustness to noise and flexibility to real-world phenomena. Figure 1 shows our adversary model based on two different assumptions related to its apriori knowledge.
@@
Figure 1: Adversary model. We consider two realistic assumptions: (i) the adversary owns a model containing features from all the devices in the pool ($\mathcal{A}_{1}$) and wants to de-anonymize any of the devices in the pool, and (ii) the adversary owns a model containing features from one only device from the pool ($\mathcal{A}_{2}$) and wants to detect its presence in the radio spectrum.
@@
Adversary 1 ($\mathcal{A}_{1}$). The adversary collects data from all devices in the pool of the transmitters, with the objective of identifying their presence in the communication link, so deanonymizing them. This objective can be formulated as a classical multiclass classification problem, where the adversary collects physical layer information from the 10 transmitters in the pool, trains a model with 10 classes (one for each transmitter), and then aims at maximizing its chances to deanonymize the transmitters (i.e., classifying them with an accuracy higher than the random guess, equal to 0.1).
@@
Adversary 2 ($\mathcal{A}_{2}$). The adversary focuses the attack on a target device, thus collecting from the physical layer of the communication link the data related to a specific transmitter. Such data is used for training a DL model to be used to detect the presence of such specific transmitter in the wild. In the remainder of this work, we address this challenge as a one-class classification problem, thus considering autoencoders as our identification tool.
4 Rationale of our Solution

### v2

# HidePrint: Protecting Device Anonymity by Obscuring Radio Fingerprints

@@ Abstract.
Radio Frequency Fingerprinting (RFF) techniques allow a receiver to authenticate a transmitter by analyzing the physical layer of the radio spectrum. Although the vast majority of scientific contributions focus on improving the performance of RFF considering different parameters and scenarios, in this work, we consider RFF as an attack vector to identify a target device in the radio spectrum.
@@
We propose, implement, and evaluate HidePrint, a solution to prevent identification through RFF without affecting the quality of the communication link between the transmitter and the receiver. HidePrint hides the transmitter’s fingerprint against an illegitimate eavesdropper through the injection of controlled noise into the transmitted signal. We evaluate our solution against various state-of-the-art RFF techniques, considering several adversarial models, data from real-world communication links (wired and wireless), and protocol configurations. Our results show that the injection of a Gaussian noise pattern with a normalized standard deviation of (at least) 0.02 prevents device fingerprinting in all the considered scenarios, while affecting the Signal-to-Noise Ratio (SNR) of the received signal by only 0.1 dB. Moreover, we introduce selective radio fingerprint disclosure, a new technique that allows the transmitter to disclose the radio fingerprint to only a subset of intended receivers.
copyright: rightsretainedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; June 1–6, 2026; Bangalore, INisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Radio Frequency Fingerprinting (RFF) techniques are today a popular tool used to differentiate and identify radio devices through the unique characteristics inherent in the Radio Frequency (RF) emissions (jagannath2022_comnet). These unique characteristics come from the hardware of the device, and in particular, from the small differences featured by the hardware components during the manufacturing process. Such differences lead to the main idea that two electronic devices that are exactly alike do not exist because of physical differences in oscillators, amplifiers, and built-in antennas. Since these differences are reflected in the over-the-air signal, a receiver can extract them through the analysis of the received signals ( In-Phase - Quadrature (IQ) samples) and leverage them to uniquely identify the transmitter (alhazbi2024_iwcmc).
A strict requirement for all RFF systems is the ability to collect information from the physical layer of the communication link (xu2016_comst). Such features persist on the receiver side up to the conversion from symbols to bits, making the availability of such information a requirement for RFF. At the physical layer, the baseband signal can be expressed via IQ samples, representing the in-phase (I) and quadrature (Q) components of the signal, used to carry the information associated with the transmitted bit string and some additional information such as noise, attenuation, Doppler shift, and finally, the radio fingerprint (shen2022_tifs).
@@
RFF techniques are characterized by several key advantages compared to traditional cryptographic solutions. RFF is crypto-less, i.e., it requires no implementation of cryptographic techniques, and, most importantly, it requires no key distribution, thus solving by design the challenges related to joining and leaving entities from the network. Moreover, RFF does not require any intervention on the transmitter side. The transmitter is authenticated through the information provided by its transmitted signal. In contrast, RFF requires a receiver capable of collecting information from the physical layer and a Deep Learning (DL) model. The receiver should train a classifier with legitimate signals (IQ samples) in order to challenge it in the subsequent phase, with a test set (of IQ samples) to assess either the presence or the absence of the transmitter(s) (danev2012_csur).
@@
Recent scientific literature mostly focused on maximizing the performance of RFF techniques while addressing several real-world challenges, e.g., robustness to multipath fading (restuccia2019_manet), jamming (smailes2024_arxiv), time-varying distortions (h2023_twc), reliability over time (alhazbi2023_acsac), and mobility of the transmitter (oligeri2023tifs), to name a few. As a result, today, RFF solutions are becoming increasingly effective for successfully classifying and authenticating the transmitter. This makes RFF a promising candidate to implement authentication at the physical layer while mitigating the common challenges of key distribution in dense networks and implementation in resource-constrained devices. While a ready-to-market RFF solution is not yet available, the accuracy of the proposed techniques is increasing for the most diverse scenarios and applications (alhazbi2024_iwcmc).
@@
In this work, we take a different perspective on RFF, exploring the possibility that a malicious party could use RFF to detect and track the presence of a device in the radio spectrum. Indeed, an adversary could train a model on signals generated by a specific (pool of) device(s) and then leverage such a model to detect the presence of the transmitter in the radio spectrum—making useless any anonymization technique implemented by higher communications layers, e.g., MAC address anonymization. Malicious detection and tracking of devices pose significant challenges to people privacy, enabling mass and targeted surveillance. For example, adversaries could track when and where a device appears in the radio spectrum to infer a person’s daily routine, places of interest, and social interactions (baig2025_tai)(ardoin2025_arxiv).
Research on protection against RFF-based attacks is relatively scarce in the literature and is mostly qualitative. Pioneering contributions consider RFF techniques for WiFi based on Channel State Information (CSI) fingerprinting (danev2010_wisec; abanto2020_macs), while more recent work explore the theoretical usage of Adversarial Machine Learning (AML) techniques to generate perturbations in the transmitted signals thwarting RFF while keeping the communication quality acceptable (papangelo2023commag)(liu2023_wcl)(sun2022_rs)(lu2024_arxiv). However, such works either focus on pure data analysis without considering the technical challenges of data manipulation and wireless transmission or use simplistic RFF models. As a result, no contributions have investigated the potential of intentionally altering the radio fingerprint to prevent the tracking of a general-purpose transmitter while allowing authorized receivers to authenticate it without affecting communication quality.
@@
RQ1: Can we prevent a malicious adversary from successfully using RFF to identify and track a (pool of) device(s)?
RQ2: Can we enable selective radio fingerprint disclosure?
Since RFF requires no shared secrets, the legitimate and malicious receivers leverage the same information. Thus, a solution preventing RFF-based attacks by an adversary (RQ1) affects the ability of a legitimate device to authenticate the transmitter(s), thus yielding to our second research question (RQ2). Selective radio fingerprint disclosure could allow a designated set of (legitimate) receivers to perform RFF while excluding other (malicious) ones.
Contribution. This work provides manifold contributions:
  • We design, implement, evaluate, and test two RFF-based attacks in different scenarios, considering both an (ideal) wired and a (real) wireless communication link. The wired setup allows us to capture the precise impact of our attacks and study in a controlled setup, while the wireless setup is used to confirm such findings in the real-world scenario.
@@
  • We consider an adversary collecting data for RFF both before and after anonymization, so updating at runtime the RFF models to de-anonymize the transmitters.
@@
  • Inspired by HidePrint, we propose a technique enabling selective radio fingerprint disclosure, thus allowing only a selected subset of legitimate receivers to perform RFF.
Paper organization. The paper is organized as follows. Sect. 2 reviews related work, Sect. 3 introduces the scenario and adversary models, Sect. 4 introduces the rationale of our solution, Sect. 5 describes our measurement campaign, Sect. 6 describes our methodology, Sect. 7 discusses the performance of our solution, Sect. 8 compares our solution to competing approaches, Sect. 9 presents selective radio fingerprint disclosure, Sect 10 discusses the main results and limitations, and finally, Sect. 11 concludes the paper.
2. Related Work
@@
Radio Fingerprint Disclosure
Technology
Independent
(danev2010_wisec) CSI-based Phase Correction Real-world tests
(abanto2020_macs) CSI-based Random Phase Correction Real-world tests
(papangelo2023commag) Image-based AML Simulations
(liu2023_wcl) Raw IQ AML Simulations
(sun2022_rs) Raw IQ AML Simulations
(lu2024_arxiv) Raw IQ AML Real-world tests
(givehchian2024_sp) CFO CFO Randomization Real-world tests
(irfan2024_asiaccs) Image-based Friendly Jamming Real-world tests
@@
Table 1. Qualitative comparison of scientific contributions focusing on RFF disruption.
@@
Early research on RFF investigated the usage of specific radio signal imperfections to identify transmitters and reject spoofing attacks, e.g., IQ imbalance, phase shift, and Carrier Frequency Offset (CFO), mostly using Machine Learning (ML) models (reising2015_tifs)(zhuang2018_asiaccs)(reising2020_iotj). With the diffusion of DL, research on RFF shifted towards using raw physical-layer signals, i.e., IQ samples, due to the enhanced capability of such algorithms to extract hidden patterns (jagannath2022_comnet). Many solutions leverage raw IQ samples directly as input to various DL classifiers, e.g., Convolutional Neural Networks (alshawabka2020_infocom) (solenthaler2025_spacesec) and Autoencoderss (smailes2023_ccs). On the one hand, such solutions are characterized by outstanding classification and authentication performance, reaching accuracy values very close to perfect discrimination (xu2016_comst). On the other hand, recent research showed that these models are very unstable, being significantly affected by any source of noise, introduced, e.g., by the warm-up of the radio (elmaghbub2024_wisec), hardware reset (alhazbi2023_acsac), temperature changes in the environment (bechir_2024_tmlcn) (gu2024_tosn), and fluctuations in channel conditions. In this context, image-based RFF techniques feature enhanced flexibility and portability across different real-world scenarios and configurations. Recent literature has shown that pre-processing (and converting) raw IQ samples into images leads to the creation of DL-based models robust to AML (papangelo2023commag), restart of the radios (alhazbi2023_acsac), and firmware reload (irfan2024_arxiv_reliability), to name a few. Image-based RFF techniques have also been used beyond device identification, e.g., for channel fingerprinting (oligeri2024sac; sadighian2024ccnc) and jamming detection (alhazbi2023ccnc; sciancalepore2023jamming).
@@
Only a few works in the literature considered RFF as an attack and, thus, investigated how to protect a device from unconditionally disclosing the radio fingerprint to malicious parties. An early attempt in this direction is the contribution by Abanto et al. (abanto2020_macs), showing that the injection of random phase corrections into the signal emitted by WiFi devices can reduce the accuracy of RFF systems using CSI to authenticate WiFi devices. The RFF algorithm considered in their paper only considers shift estimation for WiFi devices, and DL-based solutions are not considered. Also, the pioneering contribution by Danev et al. in (danev2010_wisec) considers CSI-based RFF and experiments with impersonation attacks via powerful signal generators. More recently, contributions such as the ones by Papangelo et al. (papangelo2023commag), Sun et al. (sun2022_rs), and Liu et al. (liu2023_wcl) investigated the effectiveness of AML to disrupt the correct operation of RFF. The only recent work considering the usage of adversarial techniques to conceal radio fingerprints is the preliminary work by Lu et al. in (lu2024_arxiv), providing an early solution to anonymization against RFF. The authors propose to generate a perturbation signal $\delta$ calculated according to a predefined strategy that matches the decision algorithm taken by the classifier. The adopted CNN network, designed from scratch, uses raw IQ samples, sharing the same limitations described above. Moreover, the authors use a limited set of devices (five) of different brands, easing the design of the RFF model and the following anonymization task, and use higher-order modulation schemes, increasingly sensitive to even low-power noise. Recently, Givehchian et al. (givehchian2024_sp) demonstrated that RFF methods based on the CFO can be thwarted by randomizing the CFO of Bluetooth and WiFi devices. However, their solution only works against RFF via CFO. An alternative approach is the proposal by Irfan et al. (irfan2024_asiaccs), using an external device (jammer) to protect against RFF. Although effective, such an approach requires an external device, complicating the solution design. Thus, as summarized in Tab. 1, no work in the literature investigated how to protect from malicious RFF while considering various adversary models, characterized by increasing knowledge of the RFF model of the target device(s), and state-of-the-art image-based RFF techniques, without using additional devices. Moreover, there is currently no technique providing selective radio fingerprint disclosure at the physical layer without exploiting information coming from the higher layers, e.g., channel state information (abanto2020_macs), so being independent of the adopted higher-layer technology. We address this gap in the remainder of this paper.
3. Scenario and Adversarial model
@@
Our scenario is constituted by a pool of ten (10) transmitters that aim to remain anonymous while communicating. Current state-of-the-art techniques already allow enforcing anonymity at higher layers of the ISO OSI communication stack, e.g., at the network layer via VPN and at the MAC layer via MAC randomization, while our scenario specifically considers anonymity against physical layer identification. Without loss of generality, we consider the use of the Binary Phase Shift Keying (BPSK) modulation scheme. Many modern communication technologies currently use BPSK, including IEEE 802.15.4 (Zigbee, Thread), IEEE 802.11 (WiFi), and many satellite communication systems (e.g., IRIDIUM) due to its robustness to noise introduced by the wireless channel (rappaport2024_book). Indeed, BPSK is able to handle the highest noise level (or distortions) before the demodulator takes the wrong decision on the symbol, thus representing the best modulation candidate for our objective. A transmitter willing to achieve anonymity (at the physical layer) could also switch to BPSK, thus trading off bandwidth (lower communication rate) with privacy. Moreover, we consider an adversary able to deploy state-of-the-art image-based RFF techniques to identify and (potentially) track devices in the radio spectrum. We recall that, as discussed in Sec. 2, image-based RFF models typically outperform other solutions thanks to their enhanced robustness to noise and flexibility to real-world phenomena. Figure 1 shows our adversary model based on two different assumptions related to its a priori knowledge.
@@
Figure 1. Adversary model. We consider two realistic assumptions: (i) the adversary owns a model containing features from all the devices in the pool ($\mathcal{A}_{1}$) and wants to de-anonymize any device, and (ii) the adversary owns a model containing features from one only device ($\mathcal{A}_{2}$) and wants to detect its presence in the radio spectrum.
@@
Adversary 1 ($\mathcal{A}_{1}$). The adversary collects data from all devices in the pool of the transmitters, with the objective of identifying their presence in the communication link, thus deanonymizing them. This objective can be formulated as a classical multiclass classification problem, where the adversary collects physical layer information from the 10 transmitters in the pool, trains a model with 10 classes (one for each transmitter), and then aims at maximizing its chances to deanonymize the transmitters (i.e., classifying them with an accuracy higher than the random guess, equal to 0.1).
@@
Adversary 2 ($\mathcal{A}_{2}$). The adversary focuses the attack on a target device, thus collecting the data related to a specific transmitter from the physical layer of the communication link. Such data is used for training a DL model to be used to detect the presence of such a specific transmitter in the wild. In line with the current literature (oligeri2023tifs) (smailes2023_ccs), we address this challenge as a one-class classification problem, thus considering autoencoders as our identification tool.
4. Rationale of our Solution